OpenStudy (anonymous):

A rectangular prism has a volume of nine hundred cubic centimeters, and the base has an area of fifty square centimeters. What is the height of the prism?

OpenStudy (paki):

please have a look here.... http://www.wikihow.com/Find-The-Height-Of-a-Prism

OpenStudy (nurali):

The volume of a rectangular prism is V = A h where A is the area of the base and h is the height of the prism. h=V/A h=900/50 h=18 cm

OpenStudy (anonymous):

Thank You Both. ^-^